ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 260, 2024

Aragona, R., R. Civino and N. Gavioli A modular idealizer chain and unrefinability of partitions with repeated parts

441

Basmajian, A., H. Parlier and N. G. Vlamis Bounded geometry with no bounded pants decomposition

235

Boyde, G. ℤ/p^{r}-hyperbolicity via homology  141

Brendle, T., N. Broaddus and A. Putman The high-dimensional cohomology of the moduli space of curves with level structures II: punctures and boundary

303

Broaddus, N. See T. Brendle Chanfi, D. Wonderful compactifications of Bruhat–Tits buildings in

the non-split case

105

Chong, C. T., W. Wang and Y. Yang Conservation strength of the infinite pigeonhole principle for trees

341

Civino, R. See R. Aragona

Dang, N.-T. Topological mixing of positive diagonal flows  1

Dietmann, R., A. Ostafe and I. E. Shparlinski Discriminants of fields generated by polynomials of given height

73

Gavioli, N. See R. Aragona Girão, A. and S. Letzter Immersion of complete digraphs in

Eulerian digraphs

401

Greenberg, N., J. S. Miller, A. Nies and D. Turetsky Martin– Löf reducibility and cost functions

261

Hamenstädt, U. Spotted disk and sphere graphs. II  463

Lee, M. and H. Oh Ergodic decompositions of geometric measures on Anosov homogeneous spaces

195

Letzter, S. See A. Girão Miller, J. S. See N. Greenberg Nies, A. See N. Greenberg Oh, H. See M. Lee Ostafe, A. See R. Dietmann Parlier, H. See A. Basmajian Pollicott, M. and B. Sewell An infinite interval version of the

α-Kakutani equidistribution problem

365

Putman, A. See T. Brendle Sewell, B. See M. Pollicott Shparlinski, I. E. See R. Dietmann Turetsky, D. See N. Greenberg Vlamis, N. G. See A. Basmajian Wang, W. See C. T. Chong Yang, Y. See C. T. Chong Zalesskii, P. The profinite completion of relatively hyperbolic virtu-

ally special groups 427